# On the creation of narrow AI: hierarchy and nonlocality of neural network skills

### v1

@@ 6 Related Work
Pruning. Even early approaches to pruning used second-order methods for pruning weights [54, 55], whereas our “attribution” scores are first-order. When training vision models, Zhou et al. [25] and Wen et al. [24] used a group lasso penalty like we do, albeit while training on the full data distribution. Sanh et al. [56] proposed “movement pruning” to prune weights during transfer learning. A variety of works have applied structured pruning to LLMs [57, 58], including Xia et al. [59] who develop a method for task-specific pruning. Highly relevant to our discussion here is the work of Cloud et al. [60], who apply “gradient routing” during training to localized network knowledge to different model components, allowing pruning to be used for unlearning.
Task Structure and Learning Dynamics. Several works have lately investigated the relationship between task structure and learning dynamics [12, 13, 14, 15]. Liu et al. [61] also briefly study a task similar to CMSP to show that hierarchical relationships between tasks cause what they call “domino” learning dynamics.
Machine Unlearning. Many approaches to unlearning have been proposed [62, 63, 64, 65, 66]. Guo et al. [67] study how applying unlearning fine-tuning to different model components affects unlearning success, inspired by a mechanistic understanding of how knowledge is retrieved in LLMs.
@@ 7 Discussion
Limitations: One limitation of our work is that our greedy pruning strategy is quite simple, and we cannot rule out that more sophisticated pruning strategies would be more successful in preserving some skills while unlearning others [68]. Also, we did not scale hyperparameters in our LLM experiments, and the performance of each of our models in Section 5 is likely suboptimal. We also did not evaluate the performance of our models in Section 4 and Section 5 outside their narrow distribution $\mathcal{D}_{N}$, and the paper would be more complete if we evaluated how pruning performs not only at achieving good performance on the narrow distribution, but also at unlearning skills outside that distribution.
In this work, we have studied some potential challenges involved in creating narrow AI systems, having to do both with the structure of data and the structures learned internally by neural networks. Underlying this work is a perspective from mechanistic interpretability, that neural networks compute a variety of sparse features [35], each with a distributed representation [36, 16, 46], and that these features are computed from each other hierarchically in circuits [69, 70, 71]. First, we found that in order to learn certain complex features, we may have to first train on a broad set of samples which encourage the learning of simpler features. Second, because features are distributed across model components, it is a nontrivial problem to move a set of task-specific features and circuits into a smaller network.
@@
We thank Ziming Liu, Josh Engels, David D. Baek, and Jamie Simon for helpful conversations and feedback. E.J.M. is supported by the NSF via the Graduate Research Fellowship Program (Grant No. 2141064) and under Cooperative Agreement PHY-2019786 (IAIFI).
@@ Appendix A Additional results on CMSP
Here we include some additional results on CMSP. First, in Figure 7, we provide a supplement to Figure 2 (bottom right), where here we also include loss curves of a wider network with 361 neurons instead of just 128 neurons. At this width, the single-hidden-layer networks have roughly the same total number of trainable parameters as the networks with two hidden layers of width 128. We observe that the deeper networks still learn faster than the parameter-matched shallow networks.
Figure 7: Forty runs of training with different seeds on CMSP dataset ($m=4$, $n=64$, $k=4$) $\mathcal{D}_{\{0\}}\cup\mathcal{D}_{\{1\}}\cup\mathcal{D}_{\{2\}}\cup\mathcal{D}_{\{3\}}\cup\mathcal{D}_{\{0,1,2,3\}}$. 27/40 runs converge with two hidden layers versus 19/40 runs with a single hidden layer with width 128 and 16/40 runs with a single hidden layer and width 361. More obviously however, convergence is faster with two hidden layers. Networks with two hidden layers (depth 3) and width 128 have 25602 total trainable parameters, versus networks with a single hidden layer (depth 2) have 9090 and 25633 parameters for widths 128 and 361 respectively. At width 361, the single-hidden-layer network has roughly the same parameter count as the network with two hidden layers of width 128. This experiment suggests that the beneficial effects of depth are not just due to increased network size, but due to depth.
We also show how ablation scores vary across neurons in our unregularized networks. With the setup in Section 3.3, we show ablation scores for each neuron for each subtask in Figure 8 and Figure 9. We find that there are neurons which have high scores across most subtasks, even subtasks in different “skill trees” ({0,1,2} versus {3,4,5}).
@@
Figure 9: Ablation scores for each subtask and each neuron. This is simply a different way of visualizing the results in Figure 8.
In Figure 10 we show sparsity vs accuracy curves from pruning like in Figure 3 (bottom left), across seeds and network sizes. In our unregularized networks, there is a lot of variation between runs in how entangled the different subtasks are.
Figure 10: Pruning sparsity vs accuracy curves for pretrained CMSP networks. We train MLPs of varying width and random seed on a CMSP dataset with two skill trees, as in Figure 3. We prune neurons based on ablation score on subtask {0,1,2}. On some networks, ablating neurons to maximum sparsity while preserving performance on subtask {0,1,2} robustly unlearns subtask {3,4,5}. However, for other networks, the subtasks seem more entangled, and we can recover performance on subtask {3,4,5} even after ablating to the maximum extent we can still recover 98% accuracy on subtask {0,1,2} (green line).
@@ B.1 Additional plots
We also show some additional results with LLMs. In Figure 11, we show how attribution scores correlate with ablation scores in LLMs. We show results for pruning both neurons and residual stream dimensions.
@@
In Figure 12, we show learning curves across runs when training from scratch on Python documents, when performing distillation, and when performing recovery training after pruning both neurons and residual stream dimensions, to varying target sparsities, of Llama-3.2-1B.
@@
In Figure 13, we show recovery curves after pruning neurons and residual stream dimensions from Llama-3.2-1B to varying sparsity levels. We find that pruning random components performs roughly as well as pruning components with the lowest attribution scores.
@@ B.2 Additional experimental details
When training networks from scratch and distilling from scratch (Figure 12 and Figure 6, we train transformers with the Llama 3 architecture [43] of varying size, listed in Table 1.
@@
When we prune Llama-3.2-1B, we prune neurons and residual stream dimensions with the sparsity combinations shown in Table 2.
@@ Appendix C Compute estimates
For the CMSP training experiments shown in Figure 2 and Figure 7, we ran 4 configurations each with 40 different seeds. We ran each job on a GPU, but on a cluster with a variety of different node configurations and GPUs. Jobs generally took between 5-60 minutes to complete, for between 13-160 total hours of GPU time.
For the experiments shown in Figure 3 and Figure 10, each run, across 5 choices of seed and 5 choices of width, trained networks on a CMSP task, ran pruning experiments, performed group-lasso training, and pruned again. Each such run took between 5-120 minutes on our cluster with varying GPU configurations, for a total time of 2-50 hours of GPU time.
For our MNIST experiments shown in Figure 4, we plot a total of 54 points, each of which is averaged over 10 training runs. We estimate that each run took around 2 minutes on our cluster, totaling to around 18 hours of GPU time.
For our LLM experiments, we ran our experiments on A100-80GB nodes, with a single GPU allocated per experiment. When we trained models from scratch on Python documents, we used 9 configurations with job lengths between 1-3 days. For distillation we had 9 configurations with job lengths between 2-3 days, though some jobs failed. For the group lasso training experiments in Figure 5, we show 3 configurations trained with job lengths of 3 days, and we performed recovery training on these models with a job length of 1 day. We estimate that the total time for these jobs was less than 1600 hours of A100 time, though the full set of experiments we attempted in the work that led to this manuscript could be over 5000 hours.

### v2

@@ 6 Related Work
Pruning. Even early approaches to pruning used second-order methods for pruning weights [57, 58], whereas our “attribution” scores are first-order. When training vision models, Zhou et al. [25] and Wen et al. [24] used a group lasso penalty like we do, albeit while training on the full data distribution. Sanh et al. [59] proposed “movement pruning” to prune weights during transfer learning. A variety of works have applied structured pruning to LLMs [60, 61], including Xia et al. [62] who develop a method for task-specific pruning. Vural and Erdogdu [63] provide evidence that pruning boosts sample efficiency by directing gradient updates to a sparse set of important directions.
Task Structure and Learning Dynamics. Several works have lately investigated the relationship between task structure and learning dynamics [12, 13, 14, 15]. Liu et al. [64] also briefly study a task similar to CMSP to show that hierarchical relationships between tasks cause what they call “domino” learning dynamics. We also highlight the “Skill it!” paper of Chen et al. [65], who empirically investigate hierarchical learning dependencies for real LLM tasks.
Machine Unlearning. Many approaches to unlearning have been proposed [66, 67, 68, 69, 70]. Guo et al. [71] study how applying unlearning fine-tuning to different model components affects unlearning success, inspired by a mechanistic understanding of how knowledge is retrieved in LLMs. Lee et al. [72] show that distillation can be applied after unlearning fine-tuning to more robustly remove unwanted capabilities. Highly relevant is the work of Cloud et al. [73], who apply “gradient routing” during training to localize capabilities to different model components, allowing pruning to be used for unlearning.
@@ 7 Discussion
Limitations: This work explored some aspects of neural network learning which are relevant to the creation of narrow AI systems. Much work remains to be done in determining how relevant our findings are to real-world networks in various domains. We first emphasize that CMSP is a toy task, an existence proof that curriculum learning effects can be quite strong, but it is still unclear whether these dynamics occur on more natural datasets. For our pruning experiments, we emphasize that our greedy attribution pruning approach is quite simple, and we cannot rule out that more sophisticated pruning strategies would be more successful in preserving some skills while unlearning others [74]. Also, we did not optimally scale hyperparameters in our LLM experiments, so the performance of each of our models in Section˜5 is likely imperfect. While we found that pruning a pre-existing Llama model was more efficient than training a model from scratch on Python documents, this could be due to our imperfect training setup, and not because one needs to train on a broad distribution of data to learn Python documents due to curriculum effects.
Conclusion: In this work, we have studied some potential challenges involved in creating narrow AI systems, having to do both with the structure of data and the structures learned internally by neural networks. Underlying this work is a perspective from mechanistic interpretability, that neural networks compute a variety of sparse features [35], each with a distributed representation [36, 16, 46], and that these features are computed from each other hierarchically in circuits [75, 76, 77]. First, we found that in order to learn certain complex features, we may have to first train on a broad set of samples which encourage the learning of simpler features. Second, because features are distributed across model components, it is a nontrivial problem to move a set of task-specific features and circuits into a smaller network, especially while unlearning unneeded circuits.
@@
We thank Ziming Liu, Josh Engels, David D. Baek, Tom McGrath, and Jamie Simon for helpful conversations and feedback. E.J.M. was supported by the NSF via the Graduate Research Fellowship Program (Grant No. 2141064) and under Cooperative Agreement PHY-2019786 (IAIFI).
@@ Appendix A Additional results on CMSP
Here we include some additional results on CMSP. We first explore whether network shape influences learning dynamics when learning composite tasks (w/ corresponding atomic tasks in the training dataset) vs. when learning an atomic task on its own. We find that the effect of depth is similar between composite vs. atomic tasks. This suggests that our networks do not necessarily need to be deep in order to take advantage of compositional structure in this task. Note that a previous version of this manuscript suggested the opposite–we had observed that deeper networks learn composite CMSP tasks faster than shallow networks, but it turns out that deeper networks also learn standalone atomic tasks faster than shallow networks. We show loss curves of depth-3 width-128, depth-2 width-361, and depth-2 width-128 networks in Figure˜7.
Figure 7: Impact of network shape on CMSP learning. Top: Forty runs of training with different seeds on CMSP dataset $\mathcal{D}_{\{0\}}\cup\mathcal{D}_{\{1\}}\cup\mathcal{D}_{\{2\}}\cup\mathcal{D}_{\{3\}}\cup\mathcal{D}_{\{0,1,2,3\}}$ with $m=4$, $n=64$, $k=4$. We find that 27/40 runs converge with two hidden layers versus 19/40 runs with a single hidden layer with width 128 and 16/40 runs with a single hidden layer and width 361. Single-hidden-layer networks are therefore able to learn composite subtasks in CMSP roughly as well as deeper networks. Bottom: Forty runs of training on a standalone atomic task $m=1$, $n=64$, $k=4$.
We also show how ablation scores vary across neurons in our unregularized networks. With the setup in Section˜3.3, we show ablation scores for each neuron for each subtask in Figure˜8 and Figure˜9. We find that there are neurons which have high scores across most subtasks, even subtasks in different “skill trees” ({0,1,2} versus {3,4,5}).
@@
Figure 9: Ablation scores for each subtask and each neuron. This is simply a different way of visualizing the results in Figure˜8.
In Figure˜10 we show sparsity vs accuracy curves from pruning like in Figure˜3 (bottom left), across seeds and network sizes. In our unregularized networks, there is a lot of variation between runs in how entangled the different subtasks are.
Figure 10: Pruning sparsity vs accuracy curves for pretrained CMSP networks. We train MLPs of varying width and random seed on a CMSP dataset with two skill trees, as in Figure˜3. We prune neurons based on ablation score on subtask {0,1,2}. On some networks, ablating neurons to maximum sparsity while preserving performance on subtask {0,1,2} robustly unlearns subtask {3,4,5}. However, for other networks, the subtasks seem more entangled, and we can recover performance on subtask {3,4,5} even after ablating to the maximum extent we can still recover 98% accuracy on subtask {0,1,2} (green line).
@@ B.1 Additional plots
We also show some additional results with LLMs. In Figure˜11, we show how attribution scores correlate with ablation scores in LLMs. We show results for pruning both neurons and residual stream dimensions.
@@
In Figure˜12, we show learning curves across runs when training from scratch on Python documents, when performing distillation, and when performing recovery training after pruning both neurons and residual stream dimensions, to varying target sparsities, of Llama-3.2-1B.
@@
In Figure˜13, we show recovery curves after pruning neurons and residual stream dimensions from Llama-3.2-1B to varying sparsity levels. We find that pruning random components performs roughly as well as pruning components with the lowest attribution scores.
@@ B.2 Additional experimental details
When training networks from scratch and distilling from scratch (Figure˜12 and Figure˜6, we train transformers with the Llama 3 architecture [43] of varying size, listed in Table˜2.
@@
When we prune Llama-3.2-1B, we prune neurons and residual stream dimensions with the sparsity combinations shown in Table˜3.
@@ Appendix C Compute estimates
For the CMSP training experiments shown in Figure˜2 and Figure˜7, we ran 4 configurations each with 40 different seeds. We ran each job on a GPU, but on a cluster with a variety of different node configurations and GPUs. Jobs generally took between 5-60 minutes to complete, for between 13-160 total hours of GPU time.
For the experiments shown in Figure˜3 and Figure˜10, each run, across 5 choices of seed and 5 choices of width, trained networks on a CMSP task, ran pruning experiments, performed group-lasso training, and pruned again. Each such run took between 5-120 minutes on our cluster with varying GPU configurations, for a total time of 2-50 hours of GPU time.
For our MNIST experiments shown in Figure˜4, we plot a total of 54 points, each of which is averaged over 10 training runs. We estimate that each run took around 2 minutes on our cluster, totaling to around 18 hours of GPU time.
For our LLM experiments, we ran our experiments on A100-80GB nodes, with a single GPU allocated per experiment. When we trained models from scratch on Python documents, we used 9 configurations with job lengths between 1-3 days. For distillation we had 9 configurations with job lengths between 2-3 days, though some jobs failed. For the group lasso training experiments in Figure˜5, we show 3 configurations trained with job lengths of 3 days, and we performed recovery training on these models with a job length of 1 day. We estimate that the total time for these jobs was less than 1600 hours of A100 time, though the full set of experiments we attempted in the work that led to this manuscript could be over 5000 hours.